OpenStudy (anoop27):

An electron starting from rest has a velocity v that increases linearly with time t so that v=kt, where k=2 m/s^2.Find the distance covered by it in first three seconds? a)18m b}9m c)6m d)4m

OpenStudy (anonymous):

This Can be Done in Two ways .... Do u Know Differentiation ?

OpenStudy (anonymous):

a = dv/dt = k a = 2m/s^2 S = u + 1/2 a t^2 Here u = 0 Since V= kt ( Initial t=0) so S = 1/2 * 2 * 9 = 9 m

OpenStudy (anonymous):

This can be Also Done By Using Integration S = intergral ( V) S = kt^2/2 = 2 * 9/2 = 9 m

OpenStudy (anoop27):

@Yahoo! Thanx bro(Yes I know both integration and differentiation)